# A Regret Perspective on Online Selective Generation

### v1

@@ G.4 A Proof of Theorem 2
$=\sum_{t=1}^{T}\ell_{\text{max}}^{2}\left(\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}+\sum_{\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})}\frac{p_{t}(\tau)}{\sum_{\bar{\tau}\in{\mathcal{{H}}_{t}^{C}(\tau_{t})}}p_{t}(\bar{\tau})}\right)$ (39)
$={2}\ell_{\text{max}}^{2}T,$ (40)
where (38) holds as (22), i.e., $\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}(\tau_{t})$ due to $\tau\in\mathcal{{H}}_{t}(\tau_{t})$ in the summation. Also, (39) holds due to (23), i.e., $\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}^{C}(\tau_{t})$ as $\tau\in\mathcal{{H}}_{t}^{C}(\tau_{t})$ from the summation condition.
The lower bound of potentials.
Then, we find a lower bound of $\mathbbm{E}\left[\ln\frac{W_{T}}{W_{0}}\right]$. In particular, we have
@@
$\mathbbm{E}\left[\ln{\frac{W_{T}}{W_{0}}}\right]$ $=\mathbbm{E}\left[\ln\sum_{i=1}^{|\mathcal{{H}}|}\exp\{-\eta\tilde{L}_{T}^{i}\}\right]-\ln|\mathcal{{H}}|$
@@
$\geq\mathbbm{E}\left[\ln\left(\max_{i\in\{1,\dots,|\mathcal{{H}}|\}}\exp\{-\eta\tilde{L}^{i}_{T}\}\right)\right]-\ln|\mathcal{{H}}|$
$=\mathbbm{E}\left[-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\tilde{L}^{i}_{T}\right]-\ln|\mathcal{{H}}|$
$=\mathbbm{E}\left[-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\sum_{t=1}^{T}\tilde{\ell}^{i}_{t}\right]-\ln|\mathcal{{H}}|$
$\geq-\eta\min_{i\in\{1,\dots,|\mathcal{{H}}|\}}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\tilde{\ell}^{i}_{t}-\ln|\mathcal{{H}}|$
$=-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))-\ln|\mathcal{{H}}|$
${=}-\eta\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}{\ell_{t}(\tau,\alpha)}-\ln|\mathcal{{H}}|,$
where the last equality holds as the fact of the unbiased loss estimator, i.e., $\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell(\tau_{i},\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))=\ell_{t}(\tau_{i},\alpha)$ from (29).
Combine the lower and upper bounds.
Finally, we combine the lower and upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$ to derive our regret bound. In particular, we have
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau,\alpha)$ $\leq{\eta T}\ell_{\text{max}}^{2}+\frac{\ln|\mathcal{{H}}|}{\eta},$
Then, substituting the optimal learning rate $\sqrt{\frac{\ln|\mathcal{H}|}{\ell_{\text{max}}^{2}T}}$, minimizing the upper bound, into the preceding bound yields the claimed regret guarantee, which completes the proof.
Appendix H Discussion on Selection Efficiency
@@
In this section, we demonstrate that our algorithm always converges to a non-trivial selective generator, i.e., $\operatorname*{arg\,min}_{\tau}\sum_{t=1}^{T}[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)]\notin\mathcal{{H}}^{\text{triv}}$, if there exists some $\tau\in\mathcal{{H}}\setminus\mathcal{{H}}^{\text{triv}}$ that satisfies the FDR guarantee, where $\mathcal{{H}}^{\text{triv}}\coloneqq\{\tau\mid\sum_{t=1}^{T}a_{t}(\tau)=T\}$.
@@
Let $\tau^{\text{triv}}\in\mathcal{{H}}^{\text{triv}}$, which is always abstaining, then for any $\tau^{\text{non-triv}}\in\mathcal{{H}}\setminus\mathcal{{H}}^{\text{triv}}$ that satisfies the FDR guarantee, the following inequality holds,
@@
$\frac{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right)}{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)}\leq\alpha$
@@
$\Rightarrow\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right)\leq\alpha\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right).$ (41)
@@
Let $L_{T}(\tau,\alpha)\coloneqq\sum_{t=1}^{T}[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)]$, which is the cumulative loss of $\tau$, and consider the difference of cumulative loss between $\tau^{\text{triv}}$ and $\tau^{\text{non-triv}}$,
$L_{T}(\tau^{\text{triv}})-$ $L_{T}(\tau^{\text{non-triv}})$
@@
$\qquad\qquad+\lambda\left(\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq\texttt{IDK})+\alpha\right)\bigg{]}$
@@
$\qquad\qquad+\lambda\left(\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK})+\alpha\right)\bigg{]}$
$=T+\lambda\alpha T-\Bigg{[}T-\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-non-triv}})\neq\texttt{IDK}\right)$
@@
$\qquad\qquad\qquad\qquad+\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right)\bigg{\}}\Bigg{]}$
@@
$\qquad\qquad\qquad\qquad-\lambda\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq_{E}\mathbf{y}_{t}\right).$

### v2

@@ H.4 A Proof of Theorem 1
$\mathbbm{E}_{t}\left[W_{t}\right]$ $=\mathbbm{E}_{t}\left[W_{t-1}\cdot\exp\left(\frac{2\gamma}{\ell_{\text{max}}}\left(\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))-\ell_{t}(\tau)\right)\right)\right]$
$=W_{t-1}\mathbbm{E}_{t}\left[\exp\left(\frac{2\gamma}{\ell_{\text{max}}}\left(\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))-\ell_{t}(\tau)\right)\right)\right]$
$=W_{t-1}\mathbbm{E}_{t}\left[\exp\left(\frac{2\gamma}{\ell_{\text{max}}}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\right)\right]\exp\left(-\frac{2\gamma}{\ell_{\text{max}}}\ell_{t}(\tau)\right)$
$\leq W_{t-1}\exp\left(\frac{2\gamma}{\ell_{\text{max}}}\ell_{t}(\tau)\right)\exp\left(-\frac{2\gamma}{\ell_{\text{max}}}\ell_{t}(\tau)\right)$ (68)
$=W_{t-1}.$
where (68) holds as (67). From this and $W_{0}=1$ by the definition, we have $\mathbbm{E}\left[W_{T}\right]\leq\mathbbm{E}\left[W_{T-1}\right]\leq\dots\leq 1$. By this and the Markov’s inequality, we have
@@
$\mathbbm{P}\left(\sum_{t=1}^{T}\frac{2\gamma}{\ell_{\text{max}}}\left(\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-\ell_{t}(\tau)\right)>\ln\frac{1}{\delta}\right)=\mathbbm{P}\left(W_{T}>\frac{1}{\delta}\right)\leq\delta\mathbbm{E}_{T}[W_{T}]\leq\delta.$
Then, by the union bound, for all $\tau$, we have
$\sum_{t=1}^{T}\left(\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))-\ell_{t}(\tau)\right)\leq\frac{\ell_{\text{max}}\ln(|\mathcal{{H}}|/\delta)}{2\gamma}$
with probability at least $1-\delta$, which completes the proof.
We then prove our main theorem, which consists of four steps.
First step.
We split the expected loss estimator into two logarithmic terms (Bubeck et al., 2012, Eq. (3.7)):
$\mathbbm{E}_{\tau\sim p_{t}}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))=\frac{1}{\eta_{t}}\underbrace{\ln\mathbbm{E}_{\tau\sim p_{t}}\exp\bigg(-\eta_{t}\Big(\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))-\mathbbm{E}_{\bar{\tau}\sim p_{t}}\ell_{t}(\bar{\tau}\mid\mathcal{{H}}_{t}(\tau_{t}))\Big)\bigg)}_{\text{variability term}}\\ -\frac{1}{\eta_{t}}\underbrace{\ln\mathbbm{E}_{\tau\sim p_{t}}\exp\bigg(-\eta_{t}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\bigg)}_{\text{log of an exponential expectation term}},$ (69)
where $\eta_{t}$ is non-increasing in $t\in\mathbb{N}$. In the following steps, each term will be rewritten.
Second step.
We provide an upper bound on the first term:
$\ln\mathbbm{E}_{\tau\sim p_{t}}$ $\exp\left(-\eta_{t}\Big(\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))-\mathbbm{E}_{\bar{\tau}\sim w_{t}}\ell_{t}(\bar{\tau}\mid\mathcal{{H}}_{t}(\tau_{t}))\Big)\right)$
$=\ln\mathbbm{E}_{\tau\sim p_{t}}\exp\big(-\eta_{t}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\big)+\eta_{t}\mathbbm{E}_{\bar{\tau}\sim p_{t}}\ell_{t}(\bar{\tau}\mid\mathcal{{H}}_{t}(\tau_{t}))$
$\leq\mathbbm{E}_{\tau\sim p_{t}}\left[\exp\big(-\eta_{t}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\big)-1+\eta_{t}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\right]$ (70)
$\leq\mathbbm{E}_{\tau\sim p_{t}}\eta_{t}^{2}\frac{\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))^{2}}{2}$ (71)
$\leq\frac{\eta_{t}^{2}}{2}\sum_{\tau\in\mathcal{{H}}}p_{t}(\tau)\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))^{2}$
$=\frac{\eta_{t}^{2}}{2}\sum_{\tau\in\mathcal{{H}}}p_{t}(\tau)\left(\frac{\ell_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))\right)^{2}$
$=\frac{\eta_{t}^{2}}{2}\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\tau)\left(\frac{\ell_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\right)^{2}$
$=\frac{\eta_{t}^{2}}{2}\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\tau)\left(\frac{\ell_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}\right)^{2}$ (72)
$=\frac{\eta_{t}^{2}}{2}\frac{\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\tau)\ell_{t}(\tau)^{2}}{\left(\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})\right)^{2}}$
$\leq\frac{\ell_{\text{max}}\eta_{t}^{2}}{2}\frac{\sum_{\tau\in\mathcal{{H}}_{t}(\tau_{t})}p_{t}(\tau)}{\left(\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})\right)^{2}}\ell_{t}(\tau_{t})$ (73)
$\leq\frac{\ell_{\text{max}}\eta_{t}^{2}}{2}\frac{\ell_{t}(\tau_{t})}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}p_{t}(\bar{\tau})}$
$=\frac{\ell_{\text{max}}\eta_{t}^{2}}{2}\frac{\ell_{t}(\tau)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}})),$ (74)
$=\frac{\ell_{\text{max}}\eta_{t}^{2}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t})),$ (75)
where (70) holds due to $\ln(x)\leq x-1$ for $x\leq 1$, (71) uses $\exp(x)\leq 1+x+x^{2}/2$ for $x\leq 0$, (72) holds as the same as (50), (73) holds as $\ell_{t}(\tau)=\ell_{t}(\tau^{\prime})$ for all $\tau\in\mathcal{{H}}(\tau^{\prime})$ by the definition of the specialized loss (4) and its domain $0,\ell_{\text{max}}]$, and (74) holds as the same as (52), i.e., the property of the indicator function.
Third step.
Next, we reformulate the second term, the log of an exponential expectation term. Let $\tilde{L}_{t}(\tau)\coloneqq\sum_{t=1}^{T}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))$ and $\tilde{L}_{0}(\tau)=0$. Then, we have
$-\frac{1}{\eta_{t}}\ln\mathbbm{E}_{\tau\sim p_{t}}\exp$ $\bigg(-\eta_{t}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\bigg)$
$=-\frac{1}{\eta_{t}}\ln\sum_{\tau\in\mathcal{{H}}}p_{t}(\tau)\exp\bigg(-\eta_{t}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\bigg)$
$=-\frac{1}{\eta_{t}}\ln\sum_{\tau\in\mathcal{{H}}}\frac{\exp(-\eta_{t}\tilde{L}_{t-1}(\tau))}{\sum_{\bar{\tau}\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t-1}(\bar{\tau}))}\exp\bigg(-\eta_{t}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))\bigg)$
$=-\frac{1}{\eta_{t}}\ln\frac{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t}(\tau))}{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t-1}(\tau))}.$ (76)
Fourth step.
Then, combining rewritten terms, (75) and (76), the following inequality holds:
$\sum_{t=1}^{T}\mathbbm{E}_{\tau\sim p_{t}}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))$ $\leq\sum_{t=1}^{T}\frac{\ell_{\text{max}}\eta_{t}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-\sum_{t=1}^{T}\frac{1}{\eta_{t}}\ln\frac{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t}(\tau))}{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t-1}(\tau))}$
$\leq\sum_{t=1}^{T}\frac{\ell_{\text{max}}\eta_{t}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-\frac{1}{\eta_{T}}\sum_{t=1}^{T}\ln\frac{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t}(\tau))}{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t-1}(\tau))}$ (77)
$=\sum_{t=1}^{T}\frac{\ell_{\text{max}}\eta_{t}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-\frac{1}{\eta_{T}}\ln\frac{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{T}\tilde{L}_{T}(\tau))}{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{1}\tilde{L}_{0}(\tau))}$
$=\sum_{t=1}^{T}\frac{\ell_{\text{max}}\eta_{t}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-\frac{1}{\eta_{T}}\ln\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{T}\tilde{L}_{T}(\tau))+\frac{\ln|\mathcal{{H}}|}{\eta_{T}}$
$\leq\sum_{t=1}^{T}\frac{\ell_{\text{max}}\eta_{t}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))-\frac{1}{\eta_{T}}\ln\left(\max_{\tau}\exp(-\eta_{T}\tilde{L}_{T}(\tau))\right)+\frac{\ln|\mathcal{{H}}|}{\eta_{T}}$
$=\sum_{t=1}^{T}\frac{\ell_{\text{max}}\eta_{t}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))+\min_{\tau}\sum_{t=1}^{T}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))+\frac{\ln|\mathcal{{H}}|}{\eta_{T}},$
where (77) holds as $\eta_{t}$ is non-increasing and $\ln\frac{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t}(\tau))}{\sum_{\tau\in\mathcal{{H}}}\exp(-\eta_{t}\tilde{L}_{t-1}(\tau))}\leq 0$ for all $t$.
From this and the fact (53) that $\mathbbm{E}_{\tau\sim p_{t}}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))=\ell_{t}(\tau_{t})-\gamma_{t}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))$, we have
$\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau}\sum_{t=1}^{T}\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))$ $\leq\sum_{t=1}^{T}\frac{\ell_{\text{max}}\eta_{t}}{2}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))+\frac{\ln|\mathcal{{H}}|}{\eta_{T}}+\sum_{t=1}^{T}\gamma_{t}\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))$
$=\sum_{t=1}^{T}\left(\frac{\ell_{\text{max}}\eta_{t}}{2}+\gamma_{t}\right)\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))+\frac{\ln|\mathcal{{H}}|}{\eta_{T}}.$
Then, from Lemma 54 and Corollary 63 along with the union bound, the following inequality holds with probability at least $1-2\delta^{\prime}$, where $\delta^{\prime}=\delta/2$ if $\eta_{t}\leq\frac{2\gamma_{t}}{\ell_{\text{max}}}$:
$\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau}\sum_{t=1}^{T}\ell_{t}(\tau)$ $\leq\sum_{t=1}^{T}\left(\frac{\ell_{\text{max}}\eta_{t}}{2}+\gamma_{t}\right)\ell_{t}(\tau_{t}\mid\mathcal{{H}}_{t}(\tau_{t}))+\frac{\ln|\mathcal{{H}}|}{\eta_{T}}+\frac{\ell_{\text{max}}\ln(2|\mathcal{{H}}|/\delta)}{2\gamma}$
$\leq 2\ell_{\text{max}}\sum_{t=1}^{T}\left(\frac{\ell_{\text{max}}\eta_{t}}{2}+\gamma_{t}\right)+\ell_{\text{max}}\ln{\frac{2}{\delta}}+\frac{\ln|\mathcal{{H}}|}{\eta_{T}}+\frac{\ell_{\text{max}}\ln(2|\mathcal{{H}}|/\delta)}{2\gamma},$
where the first inequality holds with probability at least $1-\delta^{\prime}$ from Corollary 63 and the second inequality holds with probability at least $1-\delta^{\prime}$ from Lemma 54 when $\alpha_{t}=\frac{\ell_{\text{max}}\eta_{t}}{2}+\gamma_{t}\leq 2\gamma_{t}$.
Final Bounds. We have our regret bound depending on the setup of the learning rate $\eta_{t}$.
If $\eta_{t}=\frac{2\gamma_{t}}{\ell_{\text{max}}}=\sqrt{\frac{\ln|\mathcal{H}|}{\ell_{\text{max}}^{2}T}}$, with probability at least $1-\delta$,
$\textbf{Reg}_{T}$ $\leq 2\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|}+\ell_{\text{max}}\left(1+\sqrt{\frac{T}{\ln|\mathcal{{H}}|}}\right)\ln\frac{2}{\delta}+2\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|}$
$=\ell_{\text{max}}\left(4\sqrt{T\ln|\mathcal{{H}}|}+\left(1+\sqrt{\frac{T}{\ln|\mathcal{{H}}|}}\right)\ln\frac{2}{\delta}\right)$
$=\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln(|\mathcal{{H}}|/\delta)}).$ (78)
If $\eta_{t}=\frac{2\gamma_{t}}{\ell_{\text{max}}}=\sqrt{\frac{\ln|\mathcal{H}|}{2\ell_{\text{max}}^{2}t}}$, with probability at least $1-\delta$,
$\textbf{Reg}_{T}$ $\leq 2\ell_{\text{max}}\sqrt{2T\ln|\mathcal{H}|}+\ell_{\text{max}}\left(1+\sqrt{\frac{2T}{\ln|\mathcal{{H}}|}}\right)\ln\frac{2}{\delta}+2\ell_{\text{max}}\sqrt{2T\ln|\mathcal{H}|}$
$=\ell_{\text{max}}\left(4\sqrt{2T\ln|\mathcal{H}|}+\left(1+\sqrt{\frac{2T}{\ln|\mathcal{{H}}|}}\right)\ln\frac{2}{\delta}\right)$
$=\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln(|\mathcal{{H}}|/\delta)}),$
noting that $\sum_{t=1}^{T}1/\sqrt{t}\leq 2\sqrt{T}$.
Expected regret
The expected regret can be obtained via integrating the deviations in (78), i.e.,
$\mathbbm{E}[W]\leq\int_{0}^{2}\frac{1}{2\delta}\mathbbm{P}\left(W>\ln\frac{2}{\delta}\right)d\delta.$
Here, if we take $W=\frac{1}{\ell_{\text{max}}\left(1+\sqrt{T/\ln|\mathcal{{H}}|}\right)}\left(\textbf{Reg}_{T}-4\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|}\right)>\ln\frac{2}{\delta}$, which holds with probability at most $\delta$, then,
$\mathbbm{E}[W]\leq 1,$
which suggests that
@@
$\mathbbm{E}[\textbf{Reg}_{T}]\leq\ell_{\text{max}}\left(4\sqrt{T\ln|\mathcal{{H}}|}+1+\sqrt{\frac{T}{\ln|\mathcal{{H}}|}}\right)=\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|}).$ (79)
Appendix I Discussion on Selection Efficiency
@@
In this section, we demonstrate that our algorithm does not converge to a trivial selective generator in the aspect of full-feedback setting, i.e., $\operatorname*{arg\,min}_{\tau}\sum_{t=1}^{T}[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)]\notin\mathcal{{H}}^{\text{triv}}$, if there exists some $\tau\in\mathcal{{H}}\setminus\mathcal{{H}}^{\text{triv}}$ that satisfies the FDR guarantee, where $\mathcal{{H}}^{\text{triv}}\coloneqq\{\tau\mid\sum_{t=1}^{T}a_{t}(\tau)=T\}$.
@@
Let $e_{t}^{\tau}\coloneqq\mathcal{{E}}_{t}(\hat{{S}}(;\tau))$, and $\tau^{\text{triv}}\in\mathcal{{H}}^{\text{triv}}$, which is always abstaining, then for any $\tau^{\text{non-triv}}\in\mathcal{{H}}\setminus\mathcal{{H}}^{\text{triv}}$ that satisfies the FDR guarantee, the following inequality holds,
@@
$\frac{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge e_{t}^{\tau^{\text{non-triv}}}\right)}{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)}\leq\alpha$
@@
$\Rightarrow\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge e_{t}^{\tau^{\text{non-triv}}}\right)\leq\alpha\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right).$ (80)
@@
Let $L_{T}(\tau,\alpha)\coloneqq\sum_{t=1}^{T}\frac{a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)}{1+\lambda}$, which is the cumulative loss of $\tau$, and consider the difference of cumulative loss between $\tau^{\text{triv}}$ and $\tau^{\text{non-triv}}$,
${(1+\lambda)}$ $(L_{T}(\tau^{\text{triv}})-L_{T}(\tau^{\text{non-triv}}))$
@@
$\qquad\qquad+\lambda\left(\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq\texttt{IDK}\wedge e_{t}^{\tau^{\text{triv}}})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{triv}})\neq\texttt{IDK})+\alpha\right)\bigg]$
@@
$\qquad\qquad+\lambda\left(\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge e_{t}^{\tau^{\text{non-triv}}})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK})+\alpha\right)\bigg]$
$=T+\lambda\alpha T-\Bigg[T-\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)$
@@
$\qquad\qquad\qquad\qquad+\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge e_{t}^{\tau^{\text{non-triv}}}\right)\bigg\}\Bigg]$
@@
$\qquad\qquad\qquad\qquad-\lambda\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\wedge e_{t}^{\tau^{\text{non-triv}}}\right).$